OpenStudy (anonymous):

derivative of 2/frthrt(x^3)

OpenStudy (anonymous):

what is frthrt?

OpenStudy (anonymous):

fourth root :))

OpenStudy (anonymous):

\[y=\frac{ 2 }{ \sqrt[4]{x^3}}=2x^{-\frac{ 3 }{ 4 }} \] and you must know that\[y=x^p \rightarrow y'=px^{p-1}\]

OpenStudy (anonymous):

\[\sqrt[b]{x^a} = x^{\frac{ a }{ b }}\]

OpenStudy (anonymous):

thankyou so much @CarlosGP and @Euler271 much help thankyou!!

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

uhmm, @CarlosGP do i have to subtract -3/4 to 1? i remember my teacher said that once you got the derivative, you have to subtract its exponent by 1..

OpenStudy (anonymous):

nope, p=-3/4 and p-1=-3/4-1=-3/4-4/4=-7/4

OpenStudy (anonymous):

okay okay, thank you so much big help for our exams.. having trouble with calculus :( thank you :)))))

OpenStudy (anonymous):

its ok